OpenStudy (noahbred):

Medal and Fan please help Simplify. | –3.9| + | 3.9| A. –7.8 B. 0 C. 7.8 D. 8

OpenStudy (hitaro9):

Are you familiar with the absolute value notation? What does it mean to take the absolute value (the vertical lines) of a number?

OpenStudy (noahbred):

its -7.8 yah

pooja195 (pooja195):

Close what is the absolute value of-3.9?

OpenStudy (hitaro9):

Can you walk me through how you got that?

OpenStudy (tkhunny):

That's a whole lot of Absolute Values to have a negative result |-3.9| = ?? |3.9| = ??

OpenStudy (noahbred):

-3.9+3.9= 7.8

OpenStudy (tkhunny):

Answer my two questions separately.

OpenStudy (noahbred):

idk how

OpenStudy (tkhunny):

Think "magnitude" only, not direction. How far is 3.9 from zero (0)?

OpenStudy (noahbred):

3 away?

OpenStudy (tkhunny):

3.9 How far from zero is -3.9?

OpenStudy (noahbred):

3.9 away

OpenStudy (tkhunny):

Perfect. You can now answer the two questions... |3.9| = ?? |-3.9| = ??

OpenStudy (noahbred):

3.9=-3.9 -3.9=3.9

OpenStudy (tkhunny):

No, |3.9| = 3.9 -- The absolute values do nothing to positive numbers. Yes, |-3.9| = 3.9 -- The absolute values change the sign of negative numbers.

OpenStudy (noahbred):

ohhh so it is positive 7.8

OpenStudy (tkhunny):

|-3.9| + |3.9| = 3.9 + 3.9 = 7.8 That's it.

OpenStudy (noahbred):

yayy thankss

OpenStudy (tkhunny):

Don't be afraid to write things out. The notation and visualization will help you. Good work.

OpenStudy (noahbred):

ok i got another one but its really different

OpenStudy (tkhunny):

Let's see it. Generally, we like one question per post. I can give you more medals that way! :-)